# Universal neighborhood topology and geometry of exceptional points in physical systems

### v1

@@ Abstract
For $N\times N$ matrices with $N>2$, the dimension of the state variable space increases with $N$, the set of linear EPs is expected to have more structure than that of the $N=2$ case. More complicated, higher-order singularities, e.g. parabolic umbilics or singularities outside of the elementary catastrophe set, may be involved, most of which have higher codimensions (minimum dimension of the parameter space to capture the full structure of the singularity set) than the elliptic umbilic. But even when a relatively complete analysis comparable to the $2\times 2$ case cannot be done, using the tools of catastrophe theory and bifurcation theory may still provide fruitful insight into the topological/geometric structure of the singularity set.
In summary, we have given a topological and geometric interpretation of EPs in linear and nonlinear physical systems, in the sense that they correspond to degenerate critical points of a (Lyapunov) potential function. Therefore, the powerful theorems of catastrophe theory can be used to identify universal features such as the shapes of the set of EPs (singularity set) as a function of the physical parameters. From this (topological) perspective, a second order linear EP can be viewed as a fourth-order nonlinear EP in an extended parameter space. Future extensions of this work may include higher-dimensional linear matrices (see, e.g., Fig. 2 in binder-kwong.2021 ) and the case where the nonlinearity couples states beyond the $2\times 2$ subset, essentially also increasing the dimension of the Hamiltonian. Besides rigorously paving the way for future design of nonlinear non-Hermitian systems with novel functionalities, the universal topology and singularity structure reported in the present study also opens up new possibilities for the targeted design of nonlinear EP sensors.
Acknowledgements.
@@
The authors gratefully acknowledge financial support for the Arizona group from the US National Science Foundation (NSF) under Grant No. DMR-1839570, and, for the Paderborn groups, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Grant No. 467358803 and the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078).

### v3

@@ 1.6 Larger matrix dimensions
For $N\times N$ matrices with $N>2$, the dimension of the state variable space increases with $N$, the set of linear EPs is expected to have more structure than that of the $N=2$ case. More complicated, higher-order singularities, e.g. parabolic umbilics or singularities outside of the elementary catastrophe set, may be involved, most of which have higher codimensions (minimum dimension of the parameter space to capture the full structure of the singularity set) than the elliptic umbilic. But even when a relatively complete analysis comparable to the $2\times 2$ case cannot be done, using the tools of catastrophe theory and bifurcation theory may still provide fruitful insight into the topological/geometric structure of the singularity set. We also note that in higher dimensional systems (described by $N\times N$ matrices with $N>2$), even in the linear case we can have EPs of various orders, for example a $3\times 3$ matrix can have second and third-order linear EPs. The structure of the nonlinear EPs can then be expected to be more complex than the case of nonlinear EPs discussed above for the $2\times 2$ matrix. We found that in the $2\times 2$ case where a Lyapunov potential exists, the coalescence of eigenvectors agrees with the condition of singularities of the Lyapunov function, and - as is already known [65, 66, 67] - the order of the nonlinear EPs can be higher than that of the linear EPs (in our case, the linear problem can only have second order EPs while the nonlinear problem has up to 4th-order nonlinear EPs).
2 Discussion
In summary, we have shown that the rich interplay of nonlinearity and non-Hermiticity at exceptional points results in a non-trivial topological structure of the EPs in the space of parameters including the nonlinearity parameter $\alpha$. This structure (or singularity set) is called elliptic umbilic and shown in Fig. 2. Our findings apply to a large class of nonlinear systems that include – but are not limited to – all power-law nonlinearity functions $f(u)$ in the effective non-Hermitian Hamiltonian Eq. (8) with self-coupling and cross-coupling, where the coupling coefficients can be real (in optics related to nonlinear refractive index coupling), imaginary (often related to nonlinear decay or gain), or complex. Asymmetric matrices with complex parameters in the off-diagonal elements also belong to this class. Further extensions of the class of matrices with the elliptic umbilic topological structure may be possible.
Specifically, we find that, by mapping EPs to degenerate critical points of a potential, catastrophe theory predicts that a second-order linear EP unfolds into the elliptic umbilic form. In practical terms, the EP perturbed by nonlinearity will spawn two touching cones of EPs with tri-cusped boundaries, independent of the underlying physical model. Interestingly, the non-trivial topological properties of the set of EPs in the space of parameters including the nonlinearity parameter $\alpha$ can be predicted solely from the Lyapunov potential at the linear EP (i.e. with $\alpha=0$). This is a strong statement made possible by catastrophe theory, and in this sense, the topology can be regarded as ‘universal’. From this topological viewpoint, the original linear EP acts as the organizing center of a higher-order singularity: a second-order EP becomes effectively fourth-order in the extended space of nonlinear parameters. These conclusions are the same for all physical platforms, e.g. photonic, atomic, and condensed-matter platforms. Our approach can be extended to the case of third (or higher) order linear EPs (e.g. [94, 95, 43, 96]) perturbed by nonlinearities. Since, in the linear case, we find the germ to be that of an elliptic umbilic catastrophe, we conjecture that all nonlinearities that are real or complex functions of the variables $w$ and $\theta$ defining the eigenvectors, and that are analytic in a neighborhood around the linear EPs, will lead to locally elliptic umbilic singularity sets.
Beyond the conceptual advance, our results have practical implications. Knowing the canonical cone–and–cusp shape of nonlinear EPs provides a clear blueprint for engineering non-Hermitian devices. For example, one can predict how tuning parameters (gain saturation, coupling strengths, etc.) will move or create EPs in polariton lasers, optical resonators, or PT-symmetric circuits. Crucially, because EP-based sensors and mode-control schemes depend on the topology of degeneracies, this universality will make it possible to establish rigorous bounds on the enhancement of sensitivity of EPs in nonlinear systems and enables the design of robust, high-sensitivity devices across different platforms. Our prediction of the singularity structure of EPs in the nonlinear regime should enable systematic studies and possible optimization of the renormalization of EPs due to noise and stochastic processes, e.g. [66, 57, 96, 34], as well as studies of Liouvillian EPs [97, 98]. From a broader perspective, identifying a common “fingerprint” for EPs highlights a deep unity in non-Hermitian physics. We expect these principles to guide future explorations – for instance in higher-dimensional multi-mode or many-body setups – and to inform the development of new spectroscopic and sensing techniques grounded in the universal topology of exceptional points.
3 Methods
The methodology of our theoretical approach is based on a branch of bifurcation theory called catastrophe theory. Here, we summarize the main aspects used in this method of analyzing nonlinear systems. More details are given in Sec. II of the Supplementary Information.
We begin with a brief review of catastrophe theory. Consider a smooth function (Lyapunov potential) $V_{L}\left(x_{1},x_{2},...,x_{n},a_{1},a_{2},...,a_{m}\right)$ of $n$ state variables $x_{i},i=1,...,n$ and $m$ control parameters $a_{j},j=1,...,m$. The critical points of $V$ are points in state space where (solutions in state space to the equations)
$\frac{\partial V_{L}}{\partial x_{i}}=0\quad,\quad i=1,...,n\,.$ (12)
Catastrophe theory seeks to classify the qualitative behavior of the critical points when the parameters are varied. If we think of the parameter space as a ’phase diagram’, qualitative changes of critical point behavior happen at parameter space points for which $V$ has degenerate critical points (singularities). At these points (in parameter and state spaces), the Hessian matrix of second derivatives is singular,
${\rm det}\left[\frac{\partial^{2}V_{L}}{\partial x_{i}\partial x_{j}}\right]=0\,$, (13)
and multiple critical points coincide. We call the set of points in parameter space where $V_{L}$ has degenerate critical points the singularity set. The theory classifies the structure of the singularity set locally around ’organizing points’. Simple structures include folds, cusps, and umbilics. Complete tables for the elementary catastrophes with codimension (defined below) up to 5 were given in Refs. [74, 75]. The table is reproduced as Supplementary Tab. I in the Supplementary Information. For each catastrophe in the table, $x\equiv x_{1}$, $y\equiv x_{2}$ are state variables, and the other lower-case letters in the Perturbation column are parameters. For a given parameterized function $V_{L}$, an isolated point in the singularity set is fixed as the organizing point (in parameter space). At this parameter value, $V_{L}$ as a function of the state variables has one or more degenerate critical points, one of which is chosen for consideration. The state coordinates are redefined with the origin shifted to this degenerate critical point, and the value of $V_{L}$ is also shifted to zero at this point. To match a particular catastrophe, one may Taylor expand $V_{L}(x,y)$, with parameters set at the organizing point values, around the origin and compare the lowest-order terms with the standard forms in the Germ column in Supplementary Tab. I. A smooth coordinate transformation may be needed to match the shown standard form. The singularity set, when restricted to the organizing point, may be structurally unstable: the geometric structure of the singularity set, and other properties, may change under small perturbations. The Perturbation column in Supplementary Tab. I. I shows some standard forms of perturbations that stabilize the singularity set: its geometric structure does not change under further small perturbations. The parameters equal zero at the organizing point. For each catastrophe, the number of parameters in the shown perturbation is the minimum number that can effect the stabilization. This minimum number is the codimension of the catastrophe. A more mathematically precise explanation of Catastrophe theory terminology and concepts, such as germs and codimension, can be found in, for example, Ref. [99].
Next, we discuss exceptional points of the (nonlinear) eigenvalue problem and the relation to catastrophe theory. Consider an eigenvalue problem with an $N\times N$ matrix which is nonlinear in the eigenvector. Suppose the eigenvector is defined by $n$ real variables (a complex variable is considered as a pair of real variables), with $n\leq 2N$. Call these variables $x_{i},i=1,...,n$. Suppose also that the matrix depends on $m$ parameters $a_{j},j=1,...,m$ and the eigenvector equation, after elimination of the eigenvalue, can be written as a set of $n$ equations
$f_{k}\left(x_{1},...,x_{n},a_{1},...,a_{m}\right)=0\quad,\quad k=1,...,n$ (14)
For each fixed parameter set $\left(a_{j},j=1,...,m\right)$, the solution $\left(x_{i},i=1,...,n\right)$ to Eq. (14) gives the eigenvector. The exceptional points are points in parameter space for which multiple eigenvectors, or equivalently, multiple solutions $\left(x_{i},i=1,...,n\right)$ to Eq. (14), coalesce.
If a potential function $V_{L}\left(x_{1},x_{2},...,x_{n},a_{1},a_{2},...,a_{m}\right)$ such that
$f_{k}\left(x_{1},...,x_{n},a_{1},...,a_{m}\right)=\frac{\partial V_{L}}{\partial x_{k}}\quad,\quad k=1,...,n$ (15)
can be constructed, then the nonlinear eigenvalue problem can be ’mapped’ onto the catastrophe theory formalism. The solutions $\left(x_{i},i=1,...,n\right)$ to Eq. (14) are the critical points of $V_{L}$, and the machinery of catastrophe theory can be used to analyze the structure of the set of exceptional points, which is just the singularity set (parameter space subset where $V_{L}$ carries degenerate critical points). In our application, the linear EP is chosen as the organizing point (see previous subsection) and structurally stabilizing perturbations are identified.
A necessary condition for the potential $V_{L}$ to exist is that the set $f_{k}$, k=1,,nk=1,...,n have ’vanishing curl’:
$\frac{\partial f_{i}}{\partial x_{j}}-\frac{\partial f_{j}}{\partial x_{i}}=0\quad,\quad i,j=1,...,n$ (16)
But for our purpose here, this condition can be slightly relaxed. Even if the $f_{k}$ in question does not satisfy Eq. (16), if a related set of functions, $f^{\prime}_{k}(x_{1},...,x_{n},a_{1},...,a_{m})$, k=1,,nk=1,...,n, has the same solutions when used in Eq. (14) and is the gradient of a potential $V_{L}$, then the critical-point analysis of $V_{L}$ is applicable to the EPs of $f_{k}$. This is the case for the eigenvector equations considered in this paper. The left hand sides of the real and imaginary parts of the eigenvector equations are equal to $w\partial V_{L}/\partial w$ and $\partial V_{L}/\partial\theta$ respectively.
4 Data availability
Source Data are provided with the paper. The main figures of this paper, Fig. 2, 3 and 4, are available as data and png files on GitHub via the DOI: https://doi.org/10.5281/zenodo.17589377
5 Code availability
The Matlab code that generated the surface of the main figure (Fig. 2), cross sections of the surface (Fig. 3), and complex eigenvalue trajectories (Fig. 4) are available on GitHub via the DOI: https://doi.org/10.5281/zenodo.17589377
Acknowledgments
@@
The authors gratefully acknowledge financial support for the Arizona group from the US National Science Foundation (NSF) under Grant No. DMR-1839570, and, for the Paderborn groups, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Grant No. 467358803 and the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078). We acknowledge support for the publication cost by the Open Access Publication Fund of Paderborn University.
Author Contributions
N.H.K, J.W., S.S. and R.B. conceived the project, N.H.K. developed the bifurcation theory, J.W. provided numerical solutions, L.A., J.S. and X.M. provided the solution algorithm for the nonlinear problem, R.B. and S.S. wrote the paper with input from all authors.
Competing interests
The authors declare no competing interests.